# Learning Contrastive Feature Representations for Facial Action Unit Detection

### v5

@@ Abstract
Existing methods for facial action unit (AU) detection often rely on encoding pixel-level information of the entire face associated with AU, which not only encodes additional redundant information but also leads to increased model complexity and limited generalizability. Additionally, the accuracy of AU detection is negatively impacted by the class imbalance issue of each AU type, and the presence of noisy and false AU labels. In this paper, we introduce a novel contrastive learning framework aimed for AU detection that incorporates both self-supervised and supervised signals, thereby enhancing the learning of discriminative features for accurate AU detection. To tackle the class imbalance issue, we employ a negative sample re-weighting strategy that adjusts the step size of updating parameters for minority and majority class samples. Moreover, to address the challenges posed by noisy and false AU labels, we employ a sampling technique that encompasses three distinct types of positive sample pairs. This enables us to inject self-supervised signals into the supervised signal, effectively mitigating the adverse effects of noisy labels. Our experimental assessments, conducted on four widely-utilized benchmark datasets (BP4D, DISFA, GFT and Aff-Wild2), underscore the superior performance of our approach compared to state-of-the-art methods of AU detection. Our code is available at https://github.com/Ziqiao-Shang/AUNCE.
@@ 1 Introduction
Facial Action Coding System (FACS) [1] is a widely employed approach for facial expression coding, which introduced a set of 44 facial action units (AUs) that establish a link between facial muscle movements and facial expressions [2]. AU detection is a multi-label binary classification problem. Each AU represents a kind of label, and activated and inactivated AUs are respectively labeled 1 and 0. Nowadays, AU detection [3, 4] has emerged as a computer vision technique to automatically identify AUs in videos or images. This advancement holds promise for diverse applications, including human-computer interaction, emotion analysis, car driving monitoring, and stuff like that.
@@
In recent years, there has been a prevailing trend in supervised frameworks for AU detection [5, 6, 7, 8, 9]. These studies have predominantly utilized the binary cross-entropy loss to strictly align predicted and ground truth labels, which is often unnecessary for classification tasks. This overemphasis on label alignment imposes excessive demands on the fitting capacity of models, consequently increasing the risk of overfitting. Moreover, as shown in Fig. 1(a), these methods commonly focus on encoding pixel-level information of the entire face associated with AU, disregarding the fact that capturing subtle feature differences when AUs activate is sufficient. As a result, these learning methods place significant demands on the model complexity to encode excessive redundant information, ultimately limiting the generalization ability of the model.
(a) Pixel-level feature learning methods
(b) Discriminative contrastive learning methods
Figure 1: Comparison between pixel-level feature learning methods and discriminative contrastive learning methods. Pixel-level feature learning approaches encode the pixel-level information of the entire face associated with AU, while discriminative contrastive learning methods encode differential information between AUs from each sample pair, thereby avoiding the encoding of redundant information irrelevant to the classification task.
@@
Representation learning approaches rooted in contrastive learning have demonstrated substantial promise [10] for classification. This approach aims to maximize the consistency of positive samples that are semantically similar while minimizing the consistency of negative samples that are semantically dissimilar [11]. By optimizing the contrastive loss, the mutual information between variables is either explicitly or implicitly optimized [12]. However, the direct integration of contrastive learning into AU detection encounters two principal challenges. The first challenge is the class imbalance issue of each AU type [7]. Most AU classes manifest a substantial imbalance between majority and minority class samples, with inactivated AUs vastly outnumbering their activated counterparts. This imbalance issue influences the classification decision threshold, potentially resulting in the direct classification of the majority of AUs as the negative class. The second challenge pertains to the noisy and false label problem [13]. Due to the subtle feature differences when AUs activate leading to the inherent difficulty in expertly labeling AUs, existing AU datasets may exhibit a few noisy and false labels, significantly undermining the overall fitting ability of models.
@@
To address these challenges, as shown in Fig. 1(b), we propose a discriminative contrastive learning framework designed to encode differential information between AUs from each sample pair, replacing the traditional pixel-level feature learning methods within the field of AU detection. The resultant loss is termed AUNCE, resembling the commonly utilized InfoNCE. Moreover, to mitigate the class imbalance issue of each AU type, we apply a negative sample re-weighting strategy, making modifications to the gradient magnitude of minority and majority class samples during back-propagation by adding importance weight to the denominator of the resultant loss AUNCE. Additionally, in response to the issue of noisy and false labeling in AUs, we introduce a positive sample samplying strategy to augment the supervision signal through self-supervised signals. This is achieved by establishing sampling ratios for various types of positive sample pairs, ensuring that more robust positive samples are derived from data augmentation with the self-supervised signal. The primary contributions of our work are succinctly summarized as follows:
@@
    We propose a discriminative contrastive learning framework to learn differential feature representations for AU detection, and the resultant loss is named AUNCE.
@@
    We employ a negative sample re-weighting strategy to tackle the class imbalance issue of each AU type.
@@
    We introduce a positive sample samplying strategy to address the issue of noisy and false AU labels.
@@
    We validate the effectiveness of AUNCE by comparing it to state-of-the-art approaches for AU detection on four widely-used datasets, namely BP4D, DISFA, GFT and Aff-Wild2.
@@ 2.1 Discriminative contrastive learning
Contrastive learning [14] initially emerged from the realm of self-supervised learning, where the self-supervised learning objectives were defined by assigning tasks to agents. Specifically, this approach encompasses the application of diverse image augmentations to identical images, followed by an assessment of the resultant images and their associated features to determine similarity. The primary objective is to augment the similarity among features extracted from the same image while concurrently attenuating the similarity between features derived from distinct images. Subsequently, the evolution of this technology has led to the extension of contrastive learning to supervised and semi-supervised learning paradigms, wherein the utilization of labeled data is leveraged to enhance the representations of model features.
@@
In recent years, discriminative contrastive learning, a subset of deep self-supervised learning methods [15], has witnessed substantial progress. This approach adheres to the learn-to-compare paradigm [16], wherein a distinction is made between observed data and noisy or erroneous data, preventing the model from engaging in the reconstruction of pixel-level information. For example, Oord et al. [12] optimized feature extraction networks by maximizing consistency between predicted and actual outcomes of sequence data, introducing the widely-utilized InfoNCE loss, which implicitly or explicitly serves as a lower bound for mutual information. Chen et al. [11] proposes a prominent framework named SimCLR, which introduces a base encoder network and a projection head trained to maximize agreement utilizing the contrastive loss InfoNCE. Additionally, MoCo, proposed by He et al. [17], employs a momentum network to maintain a queue of negative samples for efficient contrastive learning. Chen and He [18] mitigate undesired collapsing solutions, eschewing the use of negative sample pairs through a stop-gradient operation. Despite variations and errors in representation encoders (denoted as $f$) and similarity measures across different tasks [19], the underlying concept remains consistent: training the representation encoder $f$ by optimizing the contrastive loss, achieved through contrasting similar pairs ($x,x^{+}$) and dissimilar pairs ($x,x^{-}$).
@@
Apart from that, deep contrastive learning not only facilitates the automated extraction of refined feature representations through the comparison of positive and negative sample pairs but also incorporates diverse contrastive losses for optimization. Theoretical analyses concerning the generalization bounds of contrastive learning for classification tasks have been provided by Arora et al. [20]. Khosla et al. [21] extended the contrastive learning concept to supervised settings with a supervised contrastive learning loss, aiming to enhance feature expression using labeled data. Additionally, Chen et al. [22] proposed a semi-supervised contrastive learning algorithm, initially pretraining all data through contrastive learning and subsequently transferring knowledge to new models via distillation learning with the aid of labeled data.
There is a discernible shift in the construction of sample pairs for contrastive learning. In cases where data is labeled or partially labeled, positive samples can be constructed using images belonging to the same category, rather than being confined to identical images. This approach augments the diversity of positive sample pairs, consequently enhancing feature expression. Furthermore, contrastive learning serves as a precursor to supervised learning, enabling model pretraining to significantly enhance the performance of subsequent supervised learning models. However, there has been no prior work that directly applies the InfoNCE-based discriminative contrastive learning framework to AU detection task based on static images, primarily due to inherent challenges in this field: class imbalance problem of each type of AUs, noisy and erroneous annotations, and difficulty in distinguishing hard negative samples.
@@ 2.2 Contrastive learning for AU detection
Contrastive learning was firstly designed as a pretext task for unsupervised or self-supervised learning and was introduced to the domain of AU detection with limited annotations [23, 24, 25]. For example, Niu et al. [23] introduced a semi-supervised co-training approach that utilized contrastive loss to enforce conditional independence between facial representations generated by two Deep Neural Networks (DNNs). Chang et al. [24] present a knowledge-driven representation learning framework aimed for AU detection, incorporating a contrastive learning module that leverages dissimilarities among facial regions to effectively train the feature encoder. Additionally, Liu et al. [25] introduced a semi-supervised contrastive learning strategy that concurrently combined region learning, contrastive learning, and pseudo-labeling.
@@
Subsequent research endeavors extended contrastive learning to supervised settings, aiming at enhancing feature representation for AU detection through the utilization of labeled data [26, 27, 28]. Wu et al. [26] introduced a contrastive feature learning method, enabling a Convolutional Neural Network (CNN) to acquire distinctive features that represent differences between neutral and AU-activated facial images. Chen et al. [28] introduced a supervised hierarchical contrastive learning method (SupHCL) aimed at enhancing AU detection by fostering knowledge consistency across distinct facial images and various AUs. In a similar vein, Suresh et al. [27] proposed a feature-based positive matching contrastive loss, which facilitates learning distances between positives based on the similarity of corresponding AU embeddings.
However, the aforementioned methods mostly employed a combination of weighted cross-entropy loss and triplet loss to constrain the training process, which involves encoding pixel-level information of the entire face associated with AU, leading to a substantial amount of redundant information. To address this issue, Li et al. [13] formulated a self-supervised contrastive learning method, which introduced the discriminative contrastive learning framework MoCo [17] into dynamic AU detection task based on extensive amounts of unlabeled facial videos, which is aim to extract differential features between video frames, inspired by some discriminative contrastive learning techniques in video representation learning [29].
In summary, contrastive learning has demonstrated promising results in the field of AU detection. However, how to capture differential information between AUs from each sample pair instead of pixel-level information of the entire face associated with AU in AU detection task based on static images is still a problem that needs to be solved.
Figure 2: Overview of our proposed discriminative contrastive learning framework.
@@ 3 Method
In Section 3, we present a discriminative contrastive learning framework specifically designed for the task of AU detection, aiming to incentivize the encoder to learn differential information between AUs from each sample pair, rather than focusing on pixel-level information of the entire face associated with AU. Subsequently, we confront two additional challenges: 1. Tackling the issue of class imbalance pertaining to each AU type in order to enhance the accuracy of our framework. 2. Addressing the problem of noisy and false labels, aiming to bolster the robustness of our models against erroneously annotated data.
@@ 3.1 Discriminative contrastive learning framework
However, the classification methods based on cross-entropy loss suffer from a common issue of being overly constrained, which can lead to overfitting and limited generalization. This constraint becomes particularly problematic in the context of AU detection, where different AUs share a common encoder. For each AU of every sample, the predicted output $\hat{y}$ must satisfy the aforementioned condition, placing a significant burden on the fitting capability of the encoder. To address this challenge, we introduce a discriminative contrastive learning framework for AU detection. As shown in Fig. 2. It comprises a positive sample mining module that integrates self-supervised signals with supervised signals, aiming to address the issue of noisy labels resulting from subtle differences in AUs. Additionally, a negative sample weighting module is included to tackle class imbalance and hard negative sample mining challenges. These modules will be discussed in detail in Sections 3.2 and 3.3, respectively. The resultant loss is formulated as:
@@
where $f(\cdot)$ serves as an encoder, mapping the features to a $d$ dimensional vector, represents the feature representation of a specific AU. The positive sample of $x$, denoted as $x^{+}$, corresponds to samples belonging to the same class as $x$. Specifically, they are derived from three sources: samples from the same class, data augmentation of the anchor point, and feature augmentation of samples from the same class. $x^{-}$ represents negative samples of $x$, which are samples from different classes compared to $x$. $\text{sim}(f(x),f(x^{+}))$ measures the similarity between $x$ and its positive example $x^{+}$, while $\text{sim}(f(x),f(x^{-}_{j}))$ measures the similarity between the sample and its negative example. The term ${p}_{d}$ represents the data distribution, ${p}^{+}$ and ${p}^{-}$ respectively denote the class conditional distribution of positive and negative samples of $x$. Inspired by Eq (1), we introduce $w_{i}$ in Eq (2) to mitigate the influence of loss values related to commonly activated AUs in the training set. Additionally, we incorporate $\omega_{j}$ to re-weight the significance of each negative sample.
The contrastive loss ${L}_{aunce}$ distinguishes itself from weighted cross-entropy loss ${L}_{wce}$ in the following aspects: 1. The weighted cross-entropy loss is computed based on the sample and its corresponding label $y$, requiring the encoding of pixel-level information in order to align with the label. In contrast, the contrastive loss does not involve explicit labels but is computed based on the feature difference values between each pair of samples. This incentivizes the encoder to capture the differential information between AUs from each sample pair, without the need to focus on pixel-level information extraction of individual samples, reducing the borden of encoder. 2. The weighted cross-entropy loss imposes constraints on the absolute values of the predicted probabilities. However, for classification tasks, this is often unnecessary and can lead to overfitting as well as demanding higher fitting capability from the encoder. The contrastive loss focuses on constraining the similarity difference between positive and negative sample pairs without explicitly specifying their exact values. This is because the minimal value of $\mathcal{L}_{aunce}$ is 0, which is achieved when $\text{sim}((f(x),f(x^{+}))-\text{sim}(f(x),f(x^{-}_{j}))\rightarrow+\infty$ for all negative sample index $j\in{1,2,\cdots,N}$.

### v6

@@ Abstract
For the Facial Action Unit (AU) detection task, accurately capturing the subtle facial differences between distinct AUs is essential for reliable detection. Additionally, AU detection faces challenges from class imbalance and the presence of noisy or false labels, which undermine detection accuracy. In this paper, we introduce a novel contrastive learning framework aimed for AU detection that incorporates both self-supervised and supervised signals, thereby enhancing the learning of discriminative features for accurate AU detection. To tackle the class imbalance issue, we employ a negative sample re-weighting strategy that adjusts the step size of updating parameters for minority and majority class samples. Moreover, to address the challenges posed by noisy and false AU labels, we employ a sampling technique that encompasses three distinct types of positive sample pairs. This enables us to inject self-supervised signals into the supervised signal, effectively mitigating the adverse effects of noisy labels. Our experimental assessments, conducted on five widely-utilized benchmark datasets (BP4D, DISFA, BP4D+, GFT and Aff-Wild2), underscore the superior performance of our approach compared to state-of-the-art methods of AU detection. Our code is available at https://github.com/Ziqiao-Shang/AUNCE.
@@ 1 Introduction
Facial Action Coding System (FACS) [1] is a widely employed approach for facial expression coding, which introduced a set of 44 facial action units (AUs) that establish a link between facial muscle movements and facial expressions [2]. AU detection is a multi-label binary classification problem. Each AU represents a kind of label, and activated and inactivated AUs are respectively labeled 1 and 0. Nowadays, AU detection [3, 4] has emerged as a computer vision technique to automatically identify AUs in videos or images. This advancement holds promise for diverse applications, including human-computer interaction, emotion analysis, car driving monitoring et al.
@@
In recent years, supervised frameworks for AU detection have predominantly relied on binary cross-entropy loss to align predicted and ground truth labels [5, 6, 7]. While this approach is common, it often places unnecessary emphasis on strict label alignment, increasing the risk of overfitting by demanding excessive model capacity. Additionally, these methods tend to focus on encoding pixel-level facial features, overlooking the fact that capturing subtle differences in different AUs’ activation is sufficient. As a result, such approaches unnecessarily increase model complexity, limiting the generalization ability of the model.
Figure 1: Illustration of discriminative contrastive learning frameworks for representation learning. The goal is to maximize the similarity between positive samples (semantically similar) and minimize the similarity between negative samples (semantically dissimilar).
@@
As a result, representation learning approaches based on discriminative contrastive learning have emerged and shown significant potential [8] for classification tasks. As shown in Fig. 1, this approach aims to maximize the consistency of positive samples that are semantically similar while minimizing the consistency of negative samples that are semantically dissimilar. By optimizing the contrastive loss, the mutual information between variables is either explicitly or implicitly optimized [9]. However, the direct integration of contrastive learning into AU detection encounters two principal challenges. The first challenge is the class imbalance issue of each AU type [10]. Most AU classes manifest a substantial imbalance between majority and minority class samples, with inactivated AUs vastly outnumbering their activated counterparts. This imbalance issue influences the classification decision threshold, potentially resulting in the direct classification of the majority of AUs as the negative class. The second challenge pertains to the noisy and false label problem [11]. Due to the subtle feature differences when AUs activate leading to the inherent difficulty in expertly labeling AUs, existing AU datasets may exhibit a few noisy and false labels, significantly undermining the overall fitting ability of models.
@@
To address these challenges, we propose a discriminative contrastive learning framework inspired by self-supervised learning techniques for learning highly semantic image representations [8]. This framework encodes differential information between AUs from each sample pair, offering an alternative to the traditional pixel-level feature learning methods commonly employed in AU detection. The resulting loss, termed AUNCE, is conceptually similar to the widely used InfoNCE loss. Moreover, to tackle the class imbalance issue in AU detection, we adopt a negative sample re-weighting strategy inspired by [7]. This strategy addresses class imbalance by assigning varying levels of importance to individual samples during training, achieved by adjusting the weights of their loss contributions. Additionally, to mitigate the effects of noisy and false AU labels, inspired by the sample-mixing concept proposed in [12], we introduce a positive sample sampling strategy. This strategy enhances supervision signals by incorporating self-supervised signals, achieved by defining sampling ratios for different types of positive sample pairs. This ensures that more robust positive samples are derived through data augmentation combined with self-supervised signal. The primary contributions of our work are summarized as follows:
@@
    We propose a discriminative contrastive learning framework that encodes differential information between AUs from each sample pair, and the resulting loss is named AUNCE.
@@
    We adopt a negative sample re-weighting strategy to tackle the class imbalance issue of each AU type.
@@
    We introduce a positive sample samplying strategy to mitigate the effects of noisy and false AU labels.
@@
    We validate the effectiveness of AUNCE by comparing it to state-of-the-art approaches for AU detection on five widely-used datasets, namely BP4D, DISFA, BP4D+, GFT and Aff-Wild2.
@@ 2.1 Discriminative contrastive learning
Contrastive learning [9] originated in self-supervised learning, where tasks are assigned to agents by applying various augmentations to the same image and evaluating feature similarity. The goal is to maximize similarity between features of the same image while minimizing it between different images. Over time, contrastive learning has expanded to supervised and semi-supervised settings, leveraging labeled data to refine feature representations.
@@
In recent years, discriminative contrastive learning, a subset of deep self-supervised learning methods [13, 14], has witnessed substantial progress. This approach follows a learn-to-compare paradigm, distinguishing between similar (positive) and dissimilar (negative) data points in a feature space. For instance, Oord et al.[9] introduced the InfoNCE loss to optimize feature extraction by maximizing consistency in sequence data, establishing a foundation for contrastive methods. Chen et al.[13] proposed SimCLR, which uses a base encoder and projection head to maximize agreement through InfoNCE. Similarly, MoCo [14] leverages a momentum network to maintain a queue of negative samples for efficient learning. Chen and He [15] addressed collapsing solutions by avoiding negative pairs via a stop-gradient operation. While variations in representation encoders and similarity measures exist, the core principle remains: training a representation encoder $f$ by contrasting similar pairs ($x,x^{+}$) against dissimilar pairs ($x,x^{-}$) to optimize the contrastive loss.
@@
Apart from that, deep contrastive learning not only facilitates the automated extraction of refined feature representations through the comparison of positive and negative sample pairs but also incorporates diverse contrastive losses for optimization. Khosla et al. [16] extended the contrastive learning concept to supervised settings with a supervised contrastive learning loss, aiming to enhance feature expression using labeled data. Additionally, Chen et al. [17] proposed a semi-supervised contrastive learning algorithm, initially pretraining all data through contrastive learning and subsequently transferring knowledge to new models via distillation learning with the aid of labeled data.
In discriminative contrastive learning, the construction of positive sample pairs has evolved. For labeled or partially labeled data, positive pairs can be formed using images from the same category, increasing pair diversity and enhancing feature representation. Contrastive learning also supports supervised learning by enabling pretraining that boosts downstream performance. However, applying an InfoNCE-based contrastive learning framework to AU detection still faces many significant challenges, including class imbalance, noisy annotations, and the difficulty of distinguishing hard negatives.
@@ 2.2 Contrastive learning for AU detection
Contrastive learning was firstly designed as a pretext task for unsupervised or self-supervised learning and was introduced to the domain of AU detection with limited annotations [18, 19, 20]. For example, Niu et al. [18] introduced a semi-supervised co-training approach that utilized contrastive loss to enforce conditional independence between facial representations generated by two Deep Neural Networks (DNNs). Following this, Zhang et al. [19] introduced a weakly-supervised contrastive learning approach that effectively utilizes coarse-grained activity information. Additionally, Liu et al. [20] introduced a semi-supervised contrastive learning strategy that concurrently combined region learning, contrastive learning, and pseudo-labeling.
@@
Subsequent research endeavors extended contrastive learning to supervised settings, aiming at enhancing feature representation for AU detection through the utilization of labeled data [21, 22]. Chang et al. [21] present a knowledge-driven representation learning framework aimed for AU detection, incorporating a contrastive learning module that leverages dissimilarities among facial regions to effectively train the feature encoder. In a similar vein, Suresh et al. [22] proposed a feature-based positive matching contrastive loss, which facilitates learning distances between positives based on the similarity of corresponding AU embeddings.
However, previous methods primarily employed a combination of weighted cross-entropy loss and triplet loss to constrain the training process, encoding pixel-level information of the entire face associated with AUs. This results in significant redundant information that may dilute the discriminative signals necessary for accurate AU detection. To address this, Sun et al.[23] proposed EmoCo, an unsupervised discriminative contrastive learning framework that uses emotion-based image pairs for AU detection. While EmoCo improves AU detection, it faces challenges due to conflicting AU labels in emotion-based pairs, leading to false positives and false negatives, and its reliance on emotion annotations limits its applicability to datasets with only AU labels. Similarly, Li et al.[11] introduced a self-supervised discriminative contrastive learning approach using the MoCo framework [14] for dynamic AU detection in video datasets. However, this method is limited to dynamic AU detection and cannot be directly applied to static image datasets without temporal information.
In summary, while these approaches have significantly advanced the field of AU detection, each is confronted with unique challenges, including the encoding of redundant pixel-level information, reliance on emotion annotations, and limitations in handling static image datasets. Furthermore, there is currently no effective solution for applying discriminative contrastive learning methods to static image datasets that contain only AU annotations. This gap underscores the need for more adaptable and efficient strategies in AU detection.
(a) The first stage of our framework: pretraining
(b) The second stage of our framework: linear evaluation
Figure 2: Overview of the training pipeline of our discriminative contrastive learning framework. The training process consists of two stages: pretraining and linear evaluation. In the first stage, the feature encoder is pretrained by the AUNCE loss. In the second stage, the learned feature representations are utilized for AU detection, with a linear classifier trained using the weighted cross-entropy loss (WCE). The linear classifier serves to evaluate performance without heavily influencing feature learning.
@@ 3 Method
3.1 Objectives
In this study, we confront three main challenges: 1. How to design a novel framework that guides the encoder to focus on the differential information between AUs from each sample pair [23], going beyond traditional methods reliant on pixel-level facial features, 2. How to tackle the class imbalance issue across different AU types [10] to improve the overall accuracy of our framework. 3. How to handle noisy and false labels [11] to increase the model’s robustness against misannotations.
To address these challenges, Section 3 provides detailed solutions. In Section 3.2, we propose a discriminative contrastive learning framework that directs the encoder to capture differential features between AUs from each sample pair, thus moving away from conventional pixel-level feature learning. To combat the class imbalance issue, Section 3.3 introduces a negative sample re-weighting strategy, drawing inspiration from, which prioritizes underrepresented AUs and facilitates the detection of challenging samples. Besides, Section 3.4 presents a positive sample sampling strategy that incorporates augmented supervision signals, which helps the model effectively learn from noisy and false data. Finally, Section 3.5 will introduce the encoder structure we used in our framework.
@@ 3.2 Discriminative contrastive learning framework
Though cross-entropy loss is widely used and effective, it has limitations in tasks with class imbalance, complex data distributions, or fine-grained distinctions [24]. It can lead to overfitting and poor generalization, particularly without regularization or techniques like label smoothing. In AU detection, where subtle inter-class differences are critical, cross-entropy loss may restrict the model’s ability to generalize to unseen data. Additionally, the requirement for predicted outputs $\hat{y}$ to satisfy specific conditions for each AU imposes a significant burden on the encoder’s learning capacity.
To address this challenge, we introduce a discriminative contrastive learning framework for AU detection. As shown in Fig. 2, the training process of our framework consists of two stages: pretraining and linear evaluation. During the pretraining stage shown in Fig. 2(a), our framework includes not only the feature encoder but also a positive sample mining module and a negative sample re-weighting module, which together assist the training process of the encoder. The former integrates self-supervised signals with supervised signals to address the issue of noisy labels arising from subtle differences in AUs. The latter is introduced to tackle challenges related to class imbalance and hard negative sample mining. Detailed discussions of these two modules are provided in Sections 3.3 and 3.4, respectively.
The input to the encoder consists of an anchor point $x$, a positive sample $x^{+}$ that belongs to the same class as $x$, and negative samples $x^{-}$ from different classes within the same batch. These samples are passed through a shared encoder to extract feature representations. Finally, the encoder parameters are optimized using a novel contrastive loss, referred to as AUNCE, which is defined as follows:
@@
where $f(\cdot)$ serves as an encoder, mapping the features to a $d$ dimensional vector, represents the feature representation of a specific AU. $\text{sim}()$ represents the inner product similarity between two tensors. Specifically, given two tensors $\mathbf{u}$ and $\mathbf{v}$, $\text{sim}(\mathbf{u},\mathbf{v})$ is computed as their dot product, defined by $\mathbf{u}^{\top}\mathbf{v}$. $\text{sim}(f(x),f(x^{+}))$ measures the similarity between $x$ and its positive example $x^{+}$, while $\text{sim}(f(x),f(x^{-}_{j}))$ measures the similarity between the sample and its negative example. The term ${p}_{d}$ represents the data distribution, ${p}^{+}$ and ${p}^{-}$ respectively denote the class conditional distribution of positive and negative samples of $x$. Inspired by Eq (1), we introduce $w_{i}$ in Eq (2) to mitigate the influence of loss values related to commonly activated AUs in the training set. Additionally, we incorporate $\omega_{j}$ to re-weight the significance of each negative sample.
The adoption of ${L}_{aunce}$ over ${L}_{wce}$ during the pretraining phase is motivated by several key considerations: 1. The weighted cross-entropy loss is computed based on the sample and its corresponding label $y$, requiring the encoding of pixel-level information in order to align with the label. In contrast, the contrastive loss does not involve explicit labels but is computed based on the feature difference values between each pair of samples. This incentivizes the encoder to capture the differential information between AUs from each sample pair, without the need to focus on pixel-level information extraction of individual samples, reducing the borden of encoder. 2. The weighted cross-entropy loss imposes constraints on the absolute values of the predicted probabilities. However, for classification tasks, this is often unnecessary and can lead to overfitting as well as demanding higher fitting capability from the encoder. The contrastive loss focuses on constraining the similarity difference between positive and negative sample pairs without explicitly specifying their exact values. This is because the minimal value of $\mathcal{L}_{aunce}$ is 0, which is achieved when $\text{sim}((f(x),f(x^{+}))-\text{sim}(f(x),f(x^{-}_{j}))\rightarrow+\infty$ for all negative sample index $j\in{1,2,\cdots,N}$.
In evaluating the representations obtained from the pretraining process, we adhere to the established linear evaluation protocol, as commonly employed in previous studies [9, 13]. As illustrated in Fig. 2(b), a linear classifier is trained atop the encoder with fixed parameters, utilizing ${L}_{wce}$ during the training phase. It is important to note, however, that this linear classifier consists of a simple fully connected layer and serves primarily for performance evaluation rather than significantly impacting the feature learning process. Therefore, the model’s core capability is largely determined by the effectiveness of AUNCE in guiding the feature extraction.